OpenStudy (anonymous):

if the ratio of the roots of ax^2+bx+c=0 is equal to the roots of the equation x^2+x+1=0 then a,b,c are in a)A.P b)GP c)HP d)none

OpenStudy (shubhamsrg):

ratio of roots of ax^2+bx+c=0 is equal to the _______ roots of the eqn x^2+ x+1 =0 am assuming in the blank space there should be "ratio of" ?

OpenStudy (anonymous):

i didn't get u

OpenStudy (anonymous):

\[ratio= 1:x\] \[a=1\times x\] \[b=1\times x\] \[c=1\times x\]

OpenStudy (anonymous):

Isn't the answer d?

OpenStudy (anonymous):

What's HP?

OpenStudy (anonymous):

harmonic progression

OpenStudy (anonymous):

It's to do with sum and prodcut of roots.

OpenStudy (anonymous):

product*

OpenStudy (anonymous):

\[\frac{ -b }{ a }=1\] \[\frac{ c }{ a }=1\]

OpenStudy (anonymous):

The first one was meant to be -1

OpenStudy (anonymous):

not 1.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Did a typo sorry.

OpenStudy (anonymous):

it's k

OpenStudy (anonymous):

so it is not making an AP. not GP not HP....so answer is none

OpenStudy (anonymous):

Yeah I'm pretty sure it is.

OpenStudy (anonymous):

Do you have the answers? If you do, check it to see if I was correct or not.

OpenStudy (anonymous):

r u sure that statement is correct, becoz ac. to @shubhamsrg something is missing in the statement

OpenStudy (anonymous):

He thought there was a ratio given. for example. 1:2 or 3:7 etc.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

No worries.

OpenStudy (anonymous):

thank u @Azteck